# Quantum-Enhanced Neural Exchange-Correlation Functionals

### v2

@@ Abstract
Kohn-Sham Density Functional Theory (KS-DFT) provides the exact ground state energy and electron density of a molecule, contingent on the as-yet-unknown universal exchange-correlation (XC) functional. Recent research has demonstrated that neural networks can efficiently learn to represent approximations to that functional, offering accurate generalizations to molecules not present during the training process. With the latest advancements in quantum-enhanced machine learning (ML), evidence is growing that Quantum Neural Network (QNN) models may offer advantages in ML applications. In this work, we explore the use of QNNs for representing XC functionals, enhancing and comparing them to classical ML techniques. We present QNNs based on differentiable quantum circuits (DQCs) as quantum (hybrid) models for XC in KS-DFT, implemented across various architectures. We assess their performance on 1D and 3D systems. To that end, we expand existing differentiable KS-DFT frameworks and propose strategies for efficient training of such functionals, highlighting the importance of fractional orbital occupation for accurate results. Our best QNN-based XC functional yields energy profiles of the H2 and planar H4 molecules that deviate by no more than 1 mHa from the reference DMRG and FCI/6-31G results, respectively. Moreover, they reach chemical precision on a system, H2H2, not present in the training dataset, using only a few variational parameters. This work lays the foundation for the integration of quantum models in KS-DFT, thereby opening new avenues for expressing XC functionals in a differentiable way and facilitating computations of various properties.
@@ I Introduction
Density Functional Theory (DFT) has emerged as a major direction of research in modern quantum chemistry, enabling an efficient prediction of the electronic structure and properties of molecules and materials. Parr and Yang (1994); Koch and Holthausen (2015); Engel and Dreizler (2011) It has become the dominant method for solving the Schrödinger equation in physics-based atomistic simulations. Becke (2014a); Jones (2015) DFT’s effectiveness largely stems from the ability to express the ground state density of interacting electrons via a system of non-interacting electrons, the so-called Kohn-Sham system. However, DFT can yield exact ground state energy and electron density only if the exact exchange-correlation (XC) functional is provided. Despite the widespread applicability of the technique, the development of accurate and versatile DFT XC functionals remains a challenge. Cohen et al. (2012); Medvedev et al. (2017); Kepp (2017) Consequently, DFT faces problems such as underestimation of reaction barriers, band gaps Cohen et al. (2008a), and poor handling of degenerate states potentially due to delocalization and static correlation errors. Cohen et al. (2008b); Becke (2014b) The universal functional capable of solving the aforementioned problems and capturing the diverse electronic properties of molecules, solids and surfaces remains unknown. In its pursuit, hundreds of approximate XC functionals have been devised. Mardirossian and Head-Gordon (2017); Goerigk et al. (2017) Many of these XC functionals are tailored to certain types of chemical systems and properties thereof, and often require prior knowledge of the targeted data to ensure the success of the method. Caldeweyher et al. (2019); Verma and Truhlar (2020); Rohrdanz et al. (2009) An additional challenge when constructing such functionals lies in achieving a good balance between accuracy and computational efficiency. Mohr et al. (2015); Bursch et al. (2022)
@@
A number of recent studies Li et al. (2016); Brockherde et al. (2017); Nagai et al. (2020); Dick and Fernandez-Serra (2020); Kalita et al. (2021); Kirkpatrick et al. (2021); Li et al. (2021); Cuierrier et al. (2022); Riemelmoser et al. (2023) propose the use of compelling capabilities of machine learning (ML) techniques as a solution for the efficient design of XC functionals that can be used for generic DFT calculations. ML algorithms, ranging from deep neural networks to sophisticated regression models, have demonstrated their potential to unlock hidden patterns in vast datasets of chemical systems, allowing for the estimation of a universal XC functional. This combination of quantum chemistry and artificial intelligence heralds a promising era for XC functional design with many demonstrations for different systems Kasim and Vinko (2021); Nagai et al. (2022) indicating the increase of the reach of the DFT methods. The majority of these ML-based techniques employ differentiable programming as a widely used paradigm in deep learning (DL), which involves optimizing parameters using gradient-based methods during the training process. Wu et al. (2023) In particular, existing chemistry codes are being refactored to incorporate automatic differentiation. Zhang and Chan (2022); M Casares et al. (2024) Training neural XC functionals within a differentiable DFT procedure was shown to provide superior generalization capabilities in 1D Li et al. (2021); Kalita et al. (2022) compared to the previous approach. Schmidt et al. (2019a) Extensions to 3D with variation of model inputs (e.g., density and its gradients), locality Kasim and Vinko (2021); M Casares et al. (2024) and novel genetic approaches Ma et al. (2022) followed, to name just a few. Nevertheless, a universal neural XC functional has not been trained yet. Wu et al. (2023) In an effort to meet this challenge, we propose a novel method for learning representations of XC functionals.
@@
Recently, a novel DL framework within the broader scope of quantum machine learning (QML) has been introduced. Kyriienko et al. (2021) This framework utilizes differentiable quantum circuits (DQCs), capitalizing on the unique methodologies inherent in quantum computation.This category of circuits facilitates the construction of quantum neural networks (QNNs), enabling efficient function representation Abbas et al. (2021) and derivative computation through analytic differentiation rules. Mitarai et al. (2018); Schuld et al. (2019); Kyriienko and Elfving (2021) Various QNN models, which utilize different data encoding methods and distinct quantum circuit designs, have emerged. Cerezo et al. (2021a) A relevant quantum advantage in QML has been shown for situations where purely quantum data (quantum state) is used as the model input Cong et al. (2019) and when the quantum model is learning from quantum experiments. Huang et al. (2022) However, given the predominance of classical data in ML, the appeal of QNNs lies in their high effective dimension and rich expressivity, Abbas et al. (2021) which are determined by the unitary mappings of data to quantum states. Schuld et al. (2021) Despite sharing many issues with classical neural networks, such as problems with barren plateaus, Larocca et al. (2022) as well as unique issues such as the absence of efficient backpropagation, Abbas et al. (2023) advances have been made in addressing these through specific techniques Beer et al. (2020); Sack et al. (2022) and efficient output extraction methods. Huang et al. (2020); Miller et al. (2022)
Many studies have explored the combination of DFT with variational quantum algorithms through auxiliary Hamiltonians, Senjean et al. (2023) embedding schemes, Rossmannek et al. (2021) or reduced density matrix methods, Schade et al. (2022) among others. However, to the best of our knowledge, the direct application of QML for the functional representation of XC has not been explored. Therefore, in this work, we investigate the synergy between QML and DFT for designing XC functionals. pat (2024) The goal is to use techniques arising from the QML domain to process empirical data from DFT (or general quantum chemistry) calculations within the supervised ML approach to create more robust and versatile XC functionals that can be applied to a wider range of problems in the quantum chemistry domain.
@@
    Expanding of KS-DFT frameworks for quantum XC functional training.
@@
    Showcasing that quantum models can potentially be more parameter-efficient than classical ones.
This work is organized as follows: In Sec. II, we give the theoretical background of QML for KS-DFT and define our models for XC functionals. In Sec. III, we provide details of our framework implementation. In Sec. IV, we present our empirical study of the proposed architectures and discuss the required quantum ressources. Finally, in Sec. V, we discuss our conclusions and give an outlook on the development of useful quantum-enhanced XC functionals.
@@ II.1 Kohn-Sham Density Functional Theory
Contemporary KS-DFT implementations Koch and Holthausen (2015) achieve the electronic ground state through a solution of the KS equations
@@
where $\epsilon_{\mathrm{XC}}[n](\mathbf{r})$ is the XC energy per electron. Finally, the total energy is obtained as the sum, $E[n]=T_{\mathrm{S}}[n]+V[n]+E_{\mathrm{H}}[n]+E_{\mathrm{XC}}[n]$, of the kinetic energy $T_{\mathrm{S}}[n]$, the external potential energy $V[n]$, the Hartree energy $E_{\mathrm{H}}[n]$ and the XC energy $E_{\mathrm{XC}}[n]$.
@@
where $\mathbf{p}_{i}=(p^{1}_{i},...,p^{N_{b}}_{i})^{T}$, $S_{ij}=\int\mathrm{d}\mathbf{r}\xi_{i}(\mathbf{r})\xi_{j}^{*}(\mathbf{r})$ is the overlap matrix and $\mathcal{F}[n]$ is the Fock matrix as a functional of the electron density profile $n$, with elements $\mathcal{F}_{ij}=\int\mathrm{d}\mathbf{r}\,\xi_{i}(\mathbf{r})\hat{H}[n](\mathbf{r})\xi_{j}^{*}(\mathbf{r}).$ Eq. (3) is then solved, e.g., numerically by an eigendecomposition of the Fock matrix to obtain the orbital coefficients $\{\mathbf{p}_{i}\}$ that correspond to some density profile $n$. Self-consistency is achieved when the electron density no longer changes within a specified tolerance over consecutive $N_{\mathrm{KS}}$ KS iterations (i.e., solutions of Eq.(3)). The resulting energy $E$ can then be compared to validated experimental data, with the quality of the result dependent on the choice of the (neural) functional used to represent $E_{\mathrm{XC}}[n]$. If the exact XC functional was known, Hohenberg-Kohn theorems Hohenberg and Kohn (1964) guarantee that the self-consistent solution yields the exact ground-state energy and density.
@@ II.2 Embedding of neural XC functionals within KS-DFT
[…]
@@
Neural XC functionals are also characterized by various degrees of locality that represents the number of density points provided as inputs. Schmidt et al. (2019b) Locality was shown to greatly impact the ability to learn accurate XC functionals. Li et al. (2021)
@@
where the model function has access only to some $L$ neighboring density points (kernel size) at a time and is applied in succession on, e.g., $\lceil N_{\mathrm{grid}}/L\rceil$ points. Schmidt et al. (2019b) In the limit where $L=1$, the $1$-local model is represented by the functions $f^{local}_{\theta}:n_{i}\to\epsilon^{\theta}_{\text{XC},i}$. Local models are embedded in DFT via Eq. (4).
@@
$f^{\mathrm{global}}_{\theta}:\mathbb{R}^{N_{grid}}\to\mathbb{R}^{N_{grid}},\ \{n\}_{i=1}^{N_{\mathrm{grids}}}\rightarrow\{\epsilon^{\theta}_{\mathrm{XC,i}}\}_{i=1}^{N_{\mathrm{grids}}},$ (8)
@@ II.3.1 Classical architectures
The most general neural network is a Multilayer Perceptron (MLP) which can approximate any function given enough layers Hornik et al. (1989). Since we are exploring the most general quantum models, MLPs are suitable candidates for comparative studies.
@@
Finally, another notable neural network is the Global Kohn-Sham Regularizer (GKSR) model, which combines local and global convolutional layers. GKSR incorporates inductive biases such as a self-interaction gate and a negativity transform, addressing common issues like self-interaction errors and imposing the negativity condition on the XC energy. Li et al. (2021) In our implementation, the GKSR model is represented by the function:
@@ II.3.2 Quantum architectures
A widely adopted strategy for QML models involves the utilization of a QNN, typically constructed using differentiable quantum circuits. Kyriienko et al. (2021) In this work, our quantum XC functionals are based on such QNNs, which consist of three blocks: a feature map ${{U}}_{f}$, an ansatz ${{U}}_{a}$ and a measurement of an operator ${\hat{O}}$. Typically, a $N_{\mathrm{q}}$-qubit quantum computer is initialized at some initial state $|0\rangle$. Here we consider the zero computational basis state on all qubits. A feature map is used to embed the data (e.g., electron density $n$) in the Hilbert space, ${{U}}_{f}(n)|0\rangle$ in a differentiable manner for the computation of XC potential $v_{\mathrm{XC}}$. The discretized density $\{n\}_{i=1}^{N_{\mathrm{grids}}}$ values are embedded in the feature map as gate parameters. Then, an ansatz circuit ${{U}}_{a}({\theta_{a}})$ is used to act on the embedded data with local qubit operations and entangling gates that mediate interactions to represent the most general action on the data. Finally, the output of the model is extracted via projective measurements (shots) of an arbitrary cost operator $\hat{C}=\sum_{i=1}^{N_{\mathrm{o}}}\theta^{i}_{c}\hat{O}_{i}$, where $\theta_{c}$ are fixed or trainable parameters associated with Pauli operators $\hat{O}\in\{I,\hat{X},\hat{Y},\hat{Z}\}^{\otimes N_{\mathrm{q}}}$. We assume the parameters $\theta_{c}$ are non-trainable, unless specified otherwise. The cost operator is chosen to be total magnetization, $\hat{C}=\sum_{i=1}^{N_{\mathrm{q}}}\hat{Z_{i}},$ which was shown to avoid cost-function-induced barren plateaus Cerezo et al. (2021b). The representation of such model is given by a function
@@
where $U(n,{\theta})=\prod^{N_{l}}_{l=1}\left(\prod^{N_{d}}_{d=1}{{U}}^{l,d}_{a}({\theta^{l,d}_{a}}){{U}}^{l}_{f}(n)\right)$ is the general representation of the model with $N_{d}$ ansatz layers and $N_{l}$ total repetition of ansatz and feature map, which implements the data re-uploading strategy. Pérez-Salinas et al. (2020) The repetitions of layers are akin to the number of layers of classical NNs and augment the expressivity of the model. Given enough layers, classical and quantum models can, in principle, approximate a given function. Hornik et al. (1989); Pérez-Salinas et al. (2020, 2021) In Appendix C, we give details of typical building blocks of QNNs.
To parallel the local MLP, we formulate a fully Local QNN (LQNN) designed to take a single density point $n_{i}$ as input and yield the expectation of an observable:
@@
defined through Eq. (6) and integrated into DFT via Eq. (4). We explore local QNNs with both product (Eq. (16)) and Chebyshev (Eq. (17)) feature maps, named Local PrQNN and Local ChQNN, respectively. The ansatz adheres exactly to the structure specified in Eq. (27). Note that for this model it is possible to consider $L$-local case, $f^{\mathrm{LQNN}}_{\theta}:\{n_{i}\}^{L}\rightarrow\epsilon^{\theta}_{\mathrm{XC,i}}=\langle 0|U^{\dagger}(n,{\theta})\hat{C}U(n,{\theta})|0\rangle,$ where a collection of $L$ neighboring densities is embedded in a single feature map. However, as shown in ref. Li et al., 2021, only global functional (e.g., $L=N_{\mathrm{grid}}$) is capable of best approximating exact XC functionals. Hence, in the next section, we develop model architectures that embed the whole density.

### v3

@@ Abstract
Kohn-Sham Density Functional Theory (KS-DFT) provides the exact ground state energy and electron density of a molecule, contingent on the as-yet-unknown universal exchange-correlation (XC) functional. Recent research has demonstrated that neural networks can efficiently learn to represent approximations to that functional, offering accurate generalizations to molecules not present during the training process. With the latest advancements in quantum-enhanced machine learning (ML), evidence is growing that Quantum Neural Network (QNN) models may offer advantages in ML applications. In this work, we explore the use of QNNs for representing XC functionals, enhancing and comparing them to classical ML techniques. We present QNNs based on differentiable quantum circuits (DQCs) as quantum (hybrid) models for XC in KS-DFT, implemented across various architectures. We assess their performance on 1D and 3D systems. To that end, we expand existing differentiable KS-DFT frameworks and propose strategies for efficient training of such functionals, highlighting the importance of fractional orbital occupation for accurate results. We prove that, under specific assumptions, noise-induced errors do not accumulate across SCF iterations, and we confirm this result through numerical experiments. Our best QNN-based XC functional yields energy profiles of the H2 and planar H4 molecules that deviate by no more than 1 mHa from the reference DMRG and FCI/6-31G results, respectively. Moreover, they reach chemical precision on a system, H2H2, not present in the training dataset, using only a few variational parameters. This work lays the foundation for the integration of quantum models in KS-DFT, thereby opening new avenues for expressing XC functionals in a differentiable way and facilitating computations of various properties.
@@ I Introduction
Density Functional Theory (DFT) has emerged as a major direction of research in modern quantum chemistry, enabling an efficient prediction of the electronic structure and properties of molecules and materials. Parr-Yang-DFT; Koch-Holthausen-DFT; Engel-Dreizler-DFT It has become the dominant method for solving the Schrödinger equation in physics-based atomistic simulations. Becke-DFT-JCP-2014; Jones-DFT-RMP-2015 DFT’s effectiveness largely stems from the ability to express the ground state density of interacting electrons via a system of non-interacting electrons, the so-called Kohn-Sham system. However, DFT can yield exact ground state energy and electron density only if the exact exchange-correlation (XC) functional is provided. Despite the widespread applicability of the technique, the development of accurate and versatile DFT XC functionals remains a challenge. Cohen-CR-2012; Medvedev-Science-2017; Kepp-Science-2017 Consequently, DFT faces problems such as underestimation of reaction barriers, band gaps cohen2008insights, and poor handling of degenerate states potentially due to delocalization and static correlation errors. cohen2008fractional; becke2014perspective The universal functional capable of solving the aforementioned problems and capturing the diverse electronic properties of molecules, solids and surfaces remains unknown. In its pursuit, hundreds of approximate XC functionals have been devised. Mardirossian-MP-2017; Goerigk-zoo-PCCP-2017 Many of these XC functionals are tailored to certain types of chemical systems and properties thereof, and often require prior knowledge of the targeted data to ensure the success of the method. Caldeweyher-D4-JCP-2019; Verma-TC-2020; Rohrdanz-LR-JCP-2009 An additional challenge when constructing such functionals lies in achieving a good balance between accuracy and computational efficiency. Mohr-PCCP-2015; Bursch-ACIE-2022
@@
A number of studies tozer1996exchange; tozer1997calculation; Brockherde-NatCom-2017; Nagai-npj-2020; Dick-NatCom-2020; Kalita-ACR-2021; Kirkpatrick-Sci-2021; ksr; Cuierrier-JCP-2022; Riemelmoser-JCTC-2023 propose the use of compelling capabilities of machine learning (ML) techniques as a solution for the efficient design of XC functionals that can be used for generic DFT calculations. ML algorithms, ranging from deep neural networks to sophisticated regression models, have demonstrated their potential to unlock hidden patterns in vast datasets of chemical systems, allowing for the estimation of a universal XC functional. This combination of quantum chemistry and artificial intelligence heralds a promising era for XC functional design with many demonstrations for different systems kasim2021learning; nagai2022machine indicating the increase of the reach of the DFT methods. The majority of these ML-based techniques employ differentiable programming as a widely used paradigm in deep learning (DL), which involves optimizing parameters using gradient-based methods during the training process. wu2023construct In particular, existing chemistry codes are being refactored to incorporate automatic differentiation. zhang2022differentiable; casares2023graddft Training neural XC functionals within a differentiable DFT procedure was shown to provide superior generalization capabilities in 1D ksr; kalita2022well compared to the previous approach. schmidt2019machine Extensions to 3D with variation of model inputs (e.g., density and its gradients), locality kasim2021learning; casares2023graddft and novel genetic approaches ma2022evolving followed, to name just a few. Nevertheless, a universal neural XC functional has not been trained yet. wu2023construct In an effort to meet this challenge, we propose a novel method for learning representations of XC functionals.
@@
Recently, a novel DL framework within the broader scope of quantum machine learning (QML) has been introduced. kyriienko2021solving This framework utilizes differentiable quantum circuits (DQCs), capitalizing on the unique methodologies inherent in quantum computation.This category of circuits facilitates the construction of quantum neural networks (QNNs), enabling efficient function representation abbas2021power and derivative computation through analytic differentiation rules. mitarai2018quantum; schuld2019evaluating; kyriienko2021generalized Various QNN models, which utilize different data encoding methods and distinct quantum circuit designs, have emerged. cerezo2021variational A relevant quantum advantage in QML has been shown for situations where purely quantum data (quantum state) is used as the model input cong2019quantum and when the quantum model is learning from quantum experiments. huang2022quantum Given the predominance of classical data in ML, the appeal of QNNs lies in their high effective dimension and rich expressivity, abbas2021power which are determined by the unitary mappings of data to quantum states. schuld2021effect However, caution is warranted, as commonly used metrics for analyzing QNNs – such as expressivity and entangling capability – may not directly correlate with a model’s ability to perform well on a specific classical dataset. Despite sharing many issues with classical neural networks, such as problems with barren plateaus, larocca2022diagnosing as well as unique issues such as the absence of efficient backpropagation, abbas2023quantum advances have been made in addressing these through specific techniques beer2020training; sack2022avoiding and efficient output extraction methods. huang2020predicting; miller2022hardware Moreover, systematic design principles for QNN architectures that are simultaneously trainable and provably non-classically simulable have been proposed gil2024relation.
Many studies have explored the combination of DFT with variational quantum algorithms through auxiliary Hamiltonians, senjean2023toward embedding schemes, rossmannek2021quantum or reduced density matrix methods, schade2022parallel among others. However, to the best of our knowledge, the direct application of QML for the functional representation of XC has not been explored. Recent developments have further reinforced the connection between DFT and QML. Gaitan and Nori gaitan2009density established a mapping between qubit systems and fermionic lattice models, providing a foundation for employing ground-state and time-dependent DFT to study properties of large-scale quantum systems, including the determination of spectral gaps. On the other hand, Baker and Poulin baker2020density caution against overly optimistic expectations for QML, demonstrating that no QML algorithm can discover the universal functional in polynomial time unless QMA-complete problems reduce to the complexity class BQP. Within the probably approximately correct (PAC) framework of machine learning, QML can at best provide a polynomial reduction in oracle queries (training points) when learning the XC functional. We take an optimistic perspective, analogous to the situation in classical ML where theoretical arguments once predicted that training deep-learning models (e.g., due to numerous local minima) would not succeed. Nonetheless, the remarkable success of large language models has demonstrated that practical performance can exceed pessimistic theoretical limitations. Therefore, in this work, we investigate the synergy between QML and DFT for designing XC functionals. patent As suggested in ref. schuld2022quantum, our methodology incorporates quantum models as part of larger classical ML pipelines. The goal is to use techniques arising from the QML domain to process empirical data from DFT (or general quantum chemistry) calculations within the supervised ML approach to create more robust and versatile XC functionals that can be applied to a wider range of problems in the quantum chemistry domain.
@@
    Expanding of KS-DFT frameworks for quantum XC functional training and implementing in an open-source package QEX qex2025pasqal written in JAX jax2018github.
@@
    Showcasing that quantum models are at least on par with classical models in terms of accuracy and parameter-efficiency on the tested molecular systems.
  • (v)
    Providing error bounds for the KS-DFT self-consistent loop under reasonable assumptions in noisy settings, e.g., when using noisy QNN-based XC functionals.
This work is organized as follows: In Sec. II, we give the theoretical background of QML for KS-DFT and define our models for XC functionals. In Sec. III, we provide details of our framework implementation. In Sec. IV, we present our empirical study of the proposed architectures, analyse the effect of quantum and classical layers within the same backbone architecture, discuss the required quantum resources and error bounds when considering typical hardware noise. Finally, in Sec. V, we discuss our conclusions and give an outlook on the development of useful quantum-enhanced XC functionals.
@@ II.1 Kohn-Sham Density Functional Theory
Contemporary KS-DFT implementations Koch-Holthausen-DFT achieve the electronic ground state through a solution of the KS equations
@@
where $\epsilon_{\mathrm{XC}}[n](\mathbf{r})$ is the XC energy per electron. Finally, the total energy is obtained as the sum, $E[n]=T_{\mathrm{S}}[n]+V[n]+E_{\mathrm{H}}[n]+E_{\mathrm{XC}}[n]$, of the non-interacting kinetic energy $T_{\mathrm{S}}[n]$, the external potential energy $V[n]$, the Hartree energy $E_{\mathrm{H}}[n]$ and the XC energy $E_{\mathrm{XC}}[n]$.
@@
where $\mathbf{p}_{i}=(p^{1}_{i},...,p^{N_{b}}_{i})^{T}$, $S_{ij}=\int\mathrm{d}\mathbf{r}\xi_{i}(\mathbf{r})\xi_{j}^{*}(\mathbf{r})$ is the overlap matrix and $\mathcal{F}[n]$ is the Fock matrix as a functional of the electron density profile $n$, with elements $\mathcal{F}_{ij}=\int\mathrm{d}\mathbf{r}\,\xi_{i}(\mathbf{r})\hat{H}[n](\mathbf{r})\xi_{j}^{*}(\mathbf{r}).$ Eq. (3) is then solved, e.g., numerically by an eigendecomposition of the Fock matrix to obtain the orbital coefficients $\{\mathbf{p}_{i}\}$ that correspond to some density profile $n$. Self-consistency is achieved when the electron density no longer changes within a specified tolerance over consecutive $N_{\mathrm{KS}}$ KS iterations (i.e., solutions of Eq.(3)). The resulting energy $E$ can then be compared to validated experimental data, with the quality of the result dependent on the choice of the (neural) functional used to represent $E_{\mathrm{XC}}[n]$. If the exact XC functional was known, Hohenberg-Kohn theorems ks-theorem guarantee that the self-consistent solution yields the exact ground-state energy and density.
@@ II.2 Embedding of neural XC functionals within KS-DFT
In Fig. 1, we summarize the approaches to incorporate our neural XC models of different localities in KS-DFT. In particular, within the established Kohn-Sham framework, neural XC models are integrated using Eq. (2). These models are represented by differentiable functions $f_{\theta}:(n,\mathbf{r})\longrightarrow\epsilon_{\mathrm{XC}}(\mathbf{r})$ with parameters denoted by $\theta$. Consequently, the machine-learned XC energy can be expressed in general as $E_{\mathrm{XC}}[n]\approx E^{\theta}_{\mathrm{XC}}[n]=\int d\mathbf{r}f_{\theta}(n,\mathbf{r})n(\mathbf{r}),$ where $E^{\theta}_{\mathrm{XC}}[n]$ approximates the exact $E_{\mathrm{XC}}[n]$. In practice, an integration grid is used to compute the integral on a finite educated choice of $N_{\mathrm{grid}}$ grid points. The XC energy is integrated as
@@
Neural XC functionals are also characterized by various degrees of locality that represents the number of density points provided as inputs. xc-mlp-1d Locality was shown to greatly impact the ability to learn accurate XC functionals. ksr
@@
where the model function has access only to some $L$ neighboring density points (kernel size) at a time and is applied in succession on, e.g., $\lceil N_{\mathrm{grid}}/L\rceil$ points. xc-mlp-1d In the limit where $L=1$, the $1$-local model is represented by the functions $f^{local}_{\theta}:n_{i}\to\epsilon^{\theta}_{\text{XC},i}$. Local models are embedded in DFT via Eq. (4).
@@
$f^{\mathrm{global}}_{\theta}:\mathbb{R}^{N_{grid}}\to\mathbb{R}^{N_{grid}},\ \{n\}_{i=1}^{N_{\mathrm{grid}}}\rightarrow\{\epsilon^{\theta}_{\mathrm{XC,i}}\}_{i=1}^{N_{\mathrm{grid}}},$ (8)
@@ II.3.1 Classical architectures
The most general neural network is a Multilayer Perceptron (MLP) which can approximate any function given enough layers hornik1989multilayer. Since we are exploring the most general quantum models, MLPs are suitable candidates for comparative studies.
@@
Finally, another notable neural network is the Global Kohn-Sham Regularizer (GKSR) model, which combines local and global convolutional layers. GKSR incorporates inductive biases such as a self-interaction gate and a negativity transform, addressing common issues like self-interaction errors and imposing the negativity condition on the XC energy. ksr In our implementation, the GKSR model is represented by the function:
@@ II.3.2 Quantum architectures
A widely adopted strategy for QML models involves the utilization of a QNN, typically constructed using differentiable quantum circuits. kyriienko2021solving In this work, our quantum XC functionals are based on such QNNs, which consist of three blocks: a feature map ${{U}}_{f}$, an ansatz ${{U}}_{a}$ and a measurement of an operator $^{\hat{O}}$. Typically, a $N_{\mathrm{q}}$-qubit quantum computer is initialized at some initial state $|0\rangle$. Here we consider the zero computational basis state on all qubits. A feature map is used to embed the data (e.g., electron density $n$) in the Hilbert space, ${{U}}_{f}(n)|0\rangle$ in a differentiable manner for the computation of XC potential $v_{\mathrm{XC}}$. The discretized density $\{n_{i}\}_{i=1}^{N_{\mathrm{grid}}}$ values are embedded in the feature map as gate parameters. Then, an ansatz circuit ${{U}}_{a}({\theta_{a}})$ is used to act on the embedded data with local qubit operations and entangling gates that mediate interactions to represent the most general action on the data. Finally, the output of the model is extracted via projective measurements (shots) of an arbitrary cost operator $\hat{C}=\sum_{i=1}^{N_{\mathrm{o}}}\theta^{i}_{c}\hat{O}_{i}$, where $\theta_{c}$ are fixed or trainable parameters associated with Pauli operators $\hat{O}\in\{I,\hat{X},\hat{Y},\hat{Z}\}^{\otimes N_{\mathrm{q}}}$. We assume the parameters $\theta_{c}$ are non-trainable, unless specified otherwise. The cost operator is chosen to be the total magnetization, composed of local observables, $\hat{C}=\sum_{i=1}^{N_{\mathrm{q}}}\hat{Z_{i}}$, which was shown to exhibit, at worst, polynomially vanishing gradients if the depth scales as $\mathcal{O}(\text{log}(N_{q}))$, thereby mitigating cost-function-induced barren plateaus cerezo2021cost. The representation of such model is given by a function
@@
where $U(n,{\theta})=\prod^{N_{l}}_{l=1}\left(\prod^{N_{d}}_{d=1}{{U}}^{l,d}_{a}({\theta^{l,d}_{a}}){{U}}^{l}_{f}(n)\right)$ is the general representation of the model with $N_{d}$ ansatz layers and $N_{l}$ total repetition of ansatz and feature map, which implements the data re-uploading strategy. perez2020data The repetitions of layers are akin to the number of layers of classical NNs and augment the expressivity of the model. Given enough layers, classical and quantum models can, in principle, approximate a given function. hornik1989multilayer; perez2020data; perez2021one In Appendix C, we give details of typical building blocks of QNNs.
To parallel the local MLP, we formulate a fully Local QNN (LQNN) designed to take a single density point $n_{i}$, from density vector $\mathbf{n}$, as input and yield the expectation of an observable:
@@
defined through Eq. (6) and integrated into DFT via Eq. (4). We explore local QNNs with both product (Eq. (16)) and Chebyshev (Eq. (17)) feature maps, named Local PrQNN and Local ChQNN, respectively. The ansatz adheres exactly to the structure specified in Eq. (32). Note that for this model it is possible to consider $L$-local case, $f^{\mathrm{LQNN}}_{\theta}:\{n_{i}\}^{L}\rightarrow\epsilon^{\theta}_{\mathrm{XC,i}}=\langle 0|U^{\dagger}(n,{\theta})\hat{C}U(n,{\theta})|0\rangle,$ where a collection of $L$ neighboring densities is embedded in a single feature map. However, as shown in ref. ksr, only global functional (e.g., $L=N_{\mathrm{grid}}$) is capable of best approximating exact XC functionals. Hence, in the next section, we develop model architectures that embed the whole density.